OpenStudy (anonymous):

The exact mass of 1 mole of carbon-12 atoms is: a) 12.00g of pure carbon-12 b)6.02x10^23g of carbon c)6.02x10^23 carbon atoms d)6.00 amus of pure carbon-12

OpenStudy (anonymous):

ok so you are asked for mass of 1 mole of C12 now lets see how can we calculate that.. n= 1mol M=12 n=m/M --> m=n*M= 1 mol * 12 gmol-1 = 12 g